### v1

@@ 1 Introduction
Over the last two decades, sparse polynomial approximation has emerged as one of the key tools to construct surrogate models, mainly motivated by applications to uncertainty quantification; see, e.g., [5, 9, 36] and references therein. Given a high-dimensional parametric model, the idea is to approximate the parameter-to-solution map using a polynomial basis in the parametric space (cp. the concept of polynomial chaos expansion [28]). The main family of parametric models studied so far as parametric Partial Differential equations (PDEs), such as parametric elliptic equations or PDEs over parameterized domains; see, e.g., [9] and [5, Chapter 4]. These PDE models are in fact known to lead to holomorphic parameter-to-solution maps and to fast (i.e., algebraic or exponential) polynomial approximation convergence rates [10]. This has been the key ingredient behind the success of these methods (especially those based on least squares and compressed sensing), which allows them to lessen the curse of dimensionality [5]. The parametric diffusion equation is one of the most well-studied parametric PDEs in the field.
Yet, PDEs are not the only mathematical models employed to describe diffusion processes. A prominent alternative is provided by diffusion on graphs [26]. Diffusion kernels have been proposed in their discrete variant over non-Euclidean manifolds such as graph manifolds [6, 11, 26]. The information flow (i.e., the derivative of the quantity of interest with respect to time) is modeled as a weighted exchange of information between each node and its neighbors, where the rule of exchange is determined by a specific kernel. In its simplest form, the information flow is modeled by a discrete heat equation, that involves the notion of graph Laplacian [37]. Diffusion on graphs is used as a modelling tool in a variety of applications [13], such as marketing selection or influence prediction over social networks [29, 33], understanding information flowing over a network of online media sources [18], or predicting the spread of epidemics [19]. Diffusion kernels have recently gained interest by the graph deep learning community [16], leading to neural architectures in which the weights of the diffusion process are learnable [8] or a certain quantity related to a specific diffusion kernel (e.g. gradient flows [12]) is minimized.
@@ 1.1 Main contributions
    We show that the parameter-to-solution map associated with parametric graph diffusion models is holomorphic (Theorem 3.1). Leveraging this result, we prove convergence theorems for least squares and compressed sensing based polynomial surrogate models (Corollaries 2 and  3). Corollaries 2 and 3 show that it is possible to achieve quasi-optimal best $m$-term approximation rates using $m$ samples (up to log factors), hence showing that least squares and compressed sensing can lessen the curse of dimensionality for parametric graph diffusion.
@@
    We validate the proposed methodology by considering the specific case of diffusion on graphs with community structure through extensive numerical experiments. First, we consider the case of synthetic graphs generated using the stochastic block model. Then, we construct surrogate models for diffusion on a real-world graph based on Twitter data.
@@ 2.1 Graph theory: basic concepts and notation
A graph (or network) $G=(V,E)$ is a mathematical object composed of a set of nodes (or vertices) $V$ and a set of edges $E$. An edge connecting two nodes $u,v\in V$ is denoted as a couple $(u,v)\in E$. The number of nodes is denoted as $|V|$. A graph can be represented by its adjacency matrix $\boldsymbol{A}\in\mathbb{Z}^{|V|\times|V|}$, where $A_{ij}=1$ if $(i,j)\in E$, $A_{ij}=0$ otherwise. In this work we consider weighted graphs, where the adjacency matrix is replaced by a weighted adjacency matrix $\boldsymbol{W}\in\mathbb{R}^{|V|\times|V|}$. We consider networks where edge weights may change in time, i.e., $\boldsymbol{W}:[0,\infty)\rightarrow\mathbb{R}^{|V|\times|V|}$.
@@ 2.2 Diffusion on graphs
(1) $\begin{cases}\dot{\boldsymbol{u}}(t)=-c\boldsymbol{L}\boldsymbol{u}(t),&t\in[0,T]\\ \boldsymbol{u}(0)=\boldsymbol{u}_{0}&\\ \end{cases},$
where $\boldsymbol{L}=\boldsymbol{D}-\boldsymbol{A}$ is the (unweighted) graph Laplacian, $c\in\mathbb{C}$ is the (uniform) diffusion coefficient, $T>0$ is the final time and $\boldsymbol{u}_{0}$ is the initial condition. The solution, which can be derived in a few steps [30], is given by
$\boldsymbol{u}(t)=\sum_{i\in V}a_{i}(t)\boldsymbol{v}_{i},$
where $\{\boldsymbol{v}_{i}\}_{i=1}^{N}$ are the eigenvectors of $\boldsymbol{L}$ and $\{a_{i}(t)\}_{i=1}^{N}$ are the (time-dependent) coefficients of the linear expansion in the basis $\{\boldsymbol{v}_{i}\}_{i=1}^{N}$.
@@
(2) $\begin{cases}\dot{\boldsymbol{u}}(t)=\boldsymbol{M}\boldsymbol{u}(t),&t\in[0,T]\\ \boldsymbol{u}(0)=\boldsymbol{u}_{0}&\\ \end{cases},$
@@
(3) $D(\boldsymbol{C},\boldsymbol{W})_{ij}=\delta_{ij}\sum_{k\in V}C_{ik}W_{ik},\quad\forall i,j\in V,$
@@
(4) $\begin{cases}\dot{\boldsymbol{u}}(t)=\boldsymbol{M}(t)\boldsymbol{u}(t),&t\in[0,T]\\ \boldsymbol{u}(0)=\boldsymbol{u}_{0}&\\ \end{cases},$
@@
(5) $\boldsymbol{M}(t)=\boldsymbol{C}(t)\odot\boldsymbol{W}(t)-\boldsymbol{D}(\boldsymbol{C}(t),\boldsymbol{W}(t)),\quad t\in[0,T],$
@@
Problem (4) is well posed for any continuous $\boldsymbol{M}(t)$, thanks to the Picard–Lindelöf Theorem [40, Theorem 6.I] since $(t,\boldsymbol{u})\mapsto\boldsymbol{M}(t)\boldsymbol{u}$ is continuous and Lipschitz in $\boldsymbol{u}$ for any rectangular neighborhod of any arbitrary point $(\bar{t},\bar{u})\in[0,T]\times\mathbb{C}^{|V|}$. In fact, $\|\boldsymbol{M}(t)\boldsymbol{u}-\boldsymbol{M}(t)\boldsymbol{v}\|\leq\|\boldsymbol{M}(t)\|\cdot\|\boldsymbol{u}-\boldsymbol{v}\|$ and $\|\boldsymbol{M}(t)\|$ is uniformly bounded in $[0,T]$ since $\boldsymbol{M}$ is continuous.
@@ 2.3 Sparse polynomial approximation: basic concepts and methods
We proceed by introducing dome basic notions of polynomial approximation in high dimensions, mostly following the presentation in [5]. In the following, we denote $\mathbb{N}_{0}:=\{0,1,2,3,\ldots\}$. Consider a function $f:\mathcal{U}=[-1,1]^{d}\subseteq\mathbb{R}^{d}\to\mathbb{C}$, where $d>1$. Let $\varrho$ be a probability measure over $\mathcal{U}$ and $\{\psi_{\boldsymbol{\nu}}\}_{\boldsymbol{\nu}\in\mathbb{N}_{0}^{d}}$ a family of orthonormal polynomials over $L^{2}_{\varrho}(\mathcal{U})$, the space of square-summable functions on $\mathcal{U}$ with respect to the measure $\varrho$. In this setting, any function $f\in L_{\varrho}^{2}(\mathcal{U})$ has an $L^{2}_{\varrho}$-convergent orthonormal expansion
$f=\sum_{\boldsymbol{\nu}\in\mathbb{N}_{0}^{d}}c_{\boldsymbol{\nu}}\psi_{\boldsymbol{\nu}}.$
The high-dimensional polynomial basis $\{\psi_{\boldsymbol{\nu}}\}_{\boldsymbol{\nu}\in\mathbb{N}_{0}^{d}}$ is obtained as tensor product of one-dimensional orthogonal polynomials $\{\phi_{k}\}_{k\in\mathbb{N}_{0}}$ over the interval $[-1,1]$, i.e.,
$\psi_{\boldsymbol{\nu}}(\boldsymbol{x})=\prod_{k=1}^{d}\phi_{\nu_{k}}(x_{k}),\quad\boldsymbol{x}=(x_{1},\ldots,x_{d})\in\mathcal{U}.$
Furthermore, we use the multi-index notation, i.e., instead of indexing a polynomial $\psi_{i}$ by $i\in\mathbb{N}$, we index it using a multi-index $\boldsymbol{\nu}=(\nu_{1},\nu_{2},...,\nu_{d})\in\mathbb{N}_{0}^{d}$. Typical examples of $\varrho$ and $\{\psi_{\boldsymbol{\nu}}\}_{\boldsymbol{\nu}\in\mathbb{N}_{0}^{d}}$ are the uniform measure and the multivariate Legendre polynomials, respectively, defined by
(6) $\psi_{\boldsymbol{\nu}}(\boldsymbol{y})=\prod\limits_{k=1}^{d}\sqrt{2\nu_{k}+1}P_{\nu_{k}}(y_{k}),\quad\boldsymbol{y}\in[-1,1]^{d},\quad\forall\boldsymbol{\nu}\in\mathbb{N}_{0}^{d},$
where the $P_{\nu}$’s are the univariate Legendre polynomials defined, e.g., via Rodrigues’s formula:
$P_{\nu}(y)=\frac{1}{2^{\nu}\nu!}\frac{d^{\nu}}{dy^{\nu}}(y^{2}-1)^{\nu},\quad y\in[-1,1],\quad\forall\nu\in\mathbb{N}_{0}.$
Another popular choice for $\varrho$ and $\{\psi_{\boldsymbol{\nu}}\}_{\boldsymbol{\nu}\in\mathbb{N}_{0}^{d}}$ is given by the Chebyshev measure and Chebyshev polynomials, respectively; see [5] for further details.
Our goal is to find an $s$-sparse polynomial approximation to $f$. In other words, we want
$f\approx\sum_{\boldsymbol{\nu}\in S}c_{\boldsymbol{\nu}}\psi_{\boldsymbol{\nu}}.$
where $S\subseteq\mathbb{N}_{0}^{d}$ and $\lvert S\rvert\leq s$.
@@
(7) $\Lambda_{n}^{\mathrm{TD}}:=\left\{\boldsymbol{\nu}=\left(\nu_{k}\right)_{k=1}^{d}\in\mathbb{N}_{0}^{d}:\sum_{k=1}^{d}\nu_{k}\leq n\right\},$
@@
(8) $\Lambda_{n}^{\mathrm{HC}}:=\left\{\boldsymbol{\nu}=\left(\nu_{k}\right)_{k=1}^{d}\in\mathbb{N}_{0}^{d}:\prod_{k=1}^{d}\left(\nu_{k}+1\right)\leq n+1\right\}.$
Note that the parameter $n\in\mathbb{N}_{0}$, called the order of the index set, and the dimension $d$ affect the cardinality of the index set. There is a closed-form expression for the cardinality of the total degree index set [5], given by
$\lvert\Lambda_{n}^{\mathrm{TD}}\rvert=\left(\begin{array}[]{c}n+d\\ d\end{array}\right),\quad\forall d\in\mathbb{N},\quad\forall n\in\mathbb{N}_{0}.$
The cardinality of the hyperbolic cross does not have an explicit formula, but it can be estimated as
$\lvert\Lambda_{n-1}^{\mathrm{HC}}\rvert\leq\min\{2n^{3}4^{d},en^{2+\log_{2}(d)}\}$
(see [5, Appendix B] and references therein). In particular, $\lvert\Lambda_{n}^{\mathrm{HC}}\rvert\ll\lvert\Lambda_{n}^{\mathrm{TD}}\rvert$ for large $d$ or $n$. We illustrate these index sets in Figure 1, with $n=4$ and $d=3$.
(a) Total degree set of cardinality 35
(b) Hyperbolic cross set of cardinality 16
Figure 1: Examples of multi-index sets in $\mathbb{N}_{0}^{3}$ of order $n=4$.
@@
(9) $\min_{\boldsymbol{z}\in\mathbb{R}^{s}}\lVert\boldsymbol{\Psi}\boldsymbol{z}-\boldsymbol{b}\rVert_{2}^{2}.$
An approximation $\boldsymbol{\hat{c}}$ to $\boldsymbol{c}$ can be found via the Moore-Penrose pseudoinverse:
$\boldsymbol{\hat{c}}=(\boldsymbol{\Psi}^{*}\boldsymbol{\Psi})^{-1}\boldsymbol{\Psi}^{*}\boldsymbol{b}.$
In order to solve a least-squares problem, we need an overdetermined system (a system with at least as many equations as unknowns). In other words, for $\boldsymbol{\Psi}\in\mathbb{R}^{m\times s}$, we need $m\geq s$. This means that we need more sample points than coefficients $c_{\boldsymbol{\nu}_{1}},\dots,c_{\boldsymbol{\nu}_{s}}$ for this method to be effective. For further details on sparse polynomial approximation via least squares, we refer to [5, Chapter 5] and references therein.
On the other hand, suppose we have no a priori knowledge of $S$. In this case, a possible strategy is to choose a set $\Lambda\subseteq\mathbb{N}_{0}^{d}$ that contains $S$. A common choice of such $\Lambda$ in literature is represented by the hyperbolic cross, since it contains all so-called lower sets of a given cardinality; see Appendix A for further explanations). If we consider a larger set $\Lambda$ that contains $S$, we come to a point where the system is underdetermined. In other words, for $\boldsymbol{\Psi}\in\mathbb{C}^{m\times N}$, where $\Psi_{ij}=\frac{1}{\sqrt{m}}\psi_{\boldsymbol{\nu}_{j}}(\boldsymbol{x}_{i})$ for all $i=1,..,m$, $j=1,...,N$, the cardinality $N=|\Lambda|$ of the set $\Lambda$ becomes greater than the number of sample points $m$. Therefore, if we want to recover the vector $\boldsymbol{c}\in\mathbb{C}^{N}$ of coefficients, this becomes a compressed sensing problem [15], where we attempt to find the sparsest solution to the system $\boldsymbol{\Psi}\boldsymbol{c}=\boldsymbol{b}$.
@@
(10) $\min_{\boldsymbol{z}\in\mathbb{C}^{N}}\lambda\lVert\boldsymbol{z}\rVert_{1}+\lVert\boldsymbol{\Psi}\boldsymbol{z}-\boldsymbol{b}\rVert_{2}\text{, }$
@@
(11) $\min_{\boldsymbol{z}\in\mathbb{C}^{N}}\lVert\boldsymbol{z}\rVert_{1}\quad\text{ subject to }\lVert\boldsymbol{\Psi}\boldsymbol{z}-\boldsymbol{b}\rVert_{2}\leq\eta\text{, }$
@@
To enhance the compressed sensing performance, a standard approach is to replace the $\ell^{1}$-norm by a weighted $\ell^{1}$-norm. This approach relies on such weights to represent a form of prior on the unknown vector to be recovered. Given a vector of weights $\boldsymbol{w}\in\mathbb{R}^{N}$ with $\boldsymbol{w}\geq\boldsymbol{0}$, we can replace the $\ell^{1}$-norm with
$\|\boldsymbol{z}\|_{1,\boldsymbol{w}}:=\sum\limits_{i=1}^{N}w_{i}z_{i}$
in equations (10) and (11). This yields, respectively, the weighted SR-LASSO program
(12) $\min\limits_{\boldsymbol{z}\in\mathbb{C}^{N}}\lambda\|\boldsymbol{z}\|_{1,\boldsymbol{w}}+\|\boldsymbol{\Psi}\boldsymbol{z}-\boldsymbol{b}\|_{2}$
and the weighted QCBP
(13) $\min_{\boldsymbol{z}\in\mathbb{C}^{N}}\lVert\boldsymbol{z}\rVert_{1,\boldsymbol{w}}\quad\text{ subject to }\lVert\boldsymbol{\Psi}\boldsymbol{z}-\boldsymbol{b}\rVert_{2}\leq\eta.$
@@
High-dimensional functions may vary much faster in some coordinate directions than in others. This behavior is referred to as anisotropy. For such functions, choices of $S$ that treat all coordinate directions equally usually lead to poor approximations [5]. Tipically, least squares methods are preferred for known anisotropy, while compressed sensing is the standard option for unknown anisotropy. We will be using both least-squares and compressed sensing in our numerical experiments, and compare their performance in different scenarios.
@@
(14) $\sigma_{s}(\boldsymbol{c})_{p}:=\min\big{\{}\|\boldsymbol{c}-\boldsymbol{z}\|_{p}:\boldsymbol{z}\in\ell^{p}(\mathbb{N}_{0}^{d}),\;|\textnormal{supp}(\boldsymbol{z})|\leq s\big{\}}.$

### v2

@@ Abstract
journal: Computational Physics
{highlights}
Construction of sparse polynomial-based surrogate models for parametric diffusion on graphs via least squares and compressed sensing.
Derivation of convergence guarantees for least squares and compressed sensing recovery in the proposed setting.
Experimental validation on synthetic and real-world networks.
@@ 1 Introduction
Over the last two decades, sparse polynomial approximation has emerged as one of the key tools to construct surrogate models, mainly motivated by applications to uncertainty quantification; see, e.g., [1, 2, 3] and references therein. Given a high-dimensional parametric model, the idea is to approximate the parameter-to-solution map using a polynomial basis in the parametric space (cf. the concept of polynomial chaos expansion [4]). The main family of parametric models studied so far as parametric Partial Differential equations (PDEs), such as parametric elliptic equations or PDEs over parameterized domains; see, e.g., [2] and [1, Chapter 4]. These PDE models are in fact known to lead to holomorphic parameter-to-solution maps and to fast (i.e., algebraic or exponential) polynomial approximation convergence rates [5]. This has been the key ingredient behind the success of these methods (especially those based on least squares and compressed sensing), which allows them to lessen the curse of dimensionality [1]. The parametric diffusion equation is one of the most well-studied parametric PDEs in the field.
Yet, PDEs are not the only mathematical models employed to describe diffusion processes. A prominent alternative is provided by diffusion on graphs [6] (although there are well-known connections between the two areas [7, 8, 9]). Diffusion kernels have been proposed in their discrete variant over non-Euclidean manifolds such as graph manifolds [10, 11, 6]. The information flow (i.e., the derivative of the quantity of interest with respect to time) is modeled as a weighted exchange of information between each node and its neighbors, where the rule of exchange is determined by a specific kernel. In its simplest form, the information flow is modeled by a discrete heat equation, that involves the notion of graph Laplacian [12]. Diffusion on graphs is used as a modelling tool in a variety of applications [13], such as marketing selection or influence prediction over social networks [14, 15], understanding information flowing over a network of online media sources [16], or predicting the spread of epidemics [17]. Diffusion kernels have recently gained interest by the graph deep learning community [18], leading to neural architectures in which the weights of the diffusion process are learnable [19] or a certain quantity related to a specific diffusion kernel (e.g. gradient flows [20]) is minimized.
@@ 1.1 Main contributions
    We show that the parameter-to-solution map associated with parametric graph diffusion models is holomorphic (Theorem 1). Leveraging this result, we prove convergence theorems for least squares and compressed sensing based polynomial surrogate models (Corollaries 2 and  3). Corollaries 2 and 3 show that it is possible to achieve quasi-optimal best $m$-term approximation rates using $m$ samples (up to log factors), hence showing that least squares and compressed sensing can lessen the curse of dimensionality for parametric graph diffusion.
@@
    We validate the proposed methodology by considering the specific case of diffusion on graphs with community structure through extensive numerical experiments. First, we consider the case of synthetic graphs generated using the stochastic block model. Then, we construct surrogate models for diffusion on a real-world graphs based on Twitter and Facebook data, respectively.
@@ 2.1 Graph theory: basic concepts and notation
A graph (or network) $G=(V,E)$ is a mathematical object composed of a set of nodes (or vertices) $V$ and a set of edges $E$. An edge connecting two nodes $u,v\in V$ is denoted as a couple $(u,v)\in E$. The number of nodes is denoted as $|V|$. Here we focus on undirected graphs, i.e., such that $(u,v)\in E$ if and only if $(v,u)\in E$. A graph can be represented by its adjacency matrix $\boldsymbol{A}\in\{0,1\}^{|V|\times|V|}$, where $A_{ij}=1$ if $(i,j)\in E$, $A_{ij}=0$ otherwise. The degree of a node $i$ is defined as $\delta_{i}=\sum_{j=1}^{|V|}A_{ij}$, and the degree matrix is the diagonal matrix $\boldsymbol{D}\in\mathbb{N}_{0}^{|V|\times|V|}$ defined by $D_{ii}=\delta_{i}$, where $\mathbb{N}_{0}:=\{0,1,2,3,\ldots\}$. In this work we consider weighted graphs, where the adjacency matrix is replaced by a weighted adjacency matrix $\boldsymbol{W}\in\mathbb{R}^{|V|\times|V|}$. Note that the unweighted adjacency matrix (and therefore the nodes’ degree) can always be retrieved by a weighted adjacency matrix (namely, $A_{ij}=1$ if $W_{ij}\neq 0$ and $A_{ij}=0$ otherwise). We consider networks where edge weights may change in time, i.e., $\boldsymbol{W}:[0,\infty)\rightarrow\mathbb{R}^{|V|\times|V|}$.
@@
A central concept in this work is that of community. A community $\mathcal{C}$ is a subset of nodes of a graph, i.e., $\mathcal{C}\subseteq V$, defined with respect to a measure of “affinity”. For instance, in graphs generated by the Stochastic Block Model (SBM) [21], this affinity is quantified by the number of connections of every node with the other ones in the community. Communities can also be defined through a community detection algorithm, as we will see in Section 4. Communities can induce a partition on the set of nodes, although it is not a strict requirement.
@@ 2.2 Diffusion on graphs
$\begin{cases}\dot{\boldsymbol{u}}(t)=-c\boldsymbol{L}\boldsymbol{u}(t),&t\in[0,T]\\ \boldsymbol{u}(0)=\boldsymbol{u}_{0}&\\ \end{cases},$ (1)
where $\boldsymbol{L}=\boldsymbol{D}-\boldsymbol{A}$ is the (unweighted) graph Laplacian, $c\in\mathbb{C}$ is the (uniform) diffusion coefficient, $T>0$ is the final time and $\boldsymbol{u}_{0}$ is the initial condition. The solution, which can be derived in a few steps [22], is given by $\boldsymbol{u}(t)=\sum_{i\in V}a_{i}(t)\boldsymbol{v}_{i},$ where $\{\boldsymbol{v}_{i}\}_{i=1}^{N}$ are the eigenvectors of $\boldsymbol{L}$ and $\{a_{i}(t)\}_{i=1}^{N}$ are the (time-dependent) coefficients of the linear expansion in the basis $\{\boldsymbol{v}_{i}\}_{i=1}^{N}$.
@@
$\begin{cases}\dot{\boldsymbol{u}}(t)=\boldsymbol{M}\boldsymbol{u}(t),&t\in[0,T]\\ \boldsymbol{u}(0)=\boldsymbol{u}_{0}&\\ \end{cases},$ (2)
@@
$D(\boldsymbol{C},\boldsymbol{W})_{ij}=\delta_{ij}\sum_{k\in V}C_{ik}W_{ik},\quad\forall i,j\in V,$ (3)
@@
$\begin{cases}\dot{\boldsymbol{u}}(t)=\boldsymbol{M}(t)\boldsymbol{u}(t),&t\in[0,T]\\ \boldsymbol{u}(0)=\boldsymbol{u}_{0}&\\ \end{cases},$ (4)
@@
$\boldsymbol{M}(t)=\boldsymbol{C}(t)\odot\boldsymbol{W}(t)-\boldsymbol{D}(\boldsymbol{C}(t),\boldsymbol{W}(t)),\quad t\in[0,T],$ (5)
@@
Problem (4) is well posed for any continuous $\boldsymbol{M}(t)$, thanks to the Picard–Lindelöf Theorem [23, Theorem 6.I] since $(t,\boldsymbol{u})\mapsto\boldsymbol{M}(t)\boldsymbol{u}$ is continuous and Lipschitz in $\boldsymbol{u}$ for any rectangular neighborhod of any arbitrary point $(\bar{t},\bar{u})\in[0,T]\times\mathbb{C}^{|V|}$. In fact, $\|\boldsymbol{M}(t)\boldsymbol{u}-\boldsymbol{M}(t)\boldsymbol{v}\|\leq\|\boldsymbol{M}(t)\|\cdot\|\boldsymbol{u}-\boldsymbol{v}\|$ and $\|\boldsymbol{M}(t)\|$ is uniformly bounded in $[0,T]$ since $\boldsymbol{M}$ is continuous.
@@ 2.3 Sparse polynomial approximation: basic concepts and methods
We proceed by introducing some basic notions of polynomial approximation in high dimensions, mostly following the presentation in [1]. Consider a function $f:\mathcal{U}=[-1,1]^{d}\subseteq\mathbb{R}^{d}\to\mathbb{C}$, where $d>1$. Let $\varrho$ be a probability measure over $\mathcal{U}$ and $\{\psi_{\boldsymbol{\nu}}\}_{\boldsymbol{\nu}\in\mathbb{N}_{0}^{d}}$ a family of orthonormal polynomials over $L^{2}_{\varrho}(\mathcal{U})$, the space of square-summable functions on $\mathcal{U}$ with respect to the measure $\varrho$. In this setting, any function $f\in L_{\varrho}^{2}(\mathcal{U})$ has an $L^{2}_{\varrho}$-convergent orthonormal expansion $f=\sum_{\boldsymbol{\nu}\in\mathbb{N}_{0}^{d}}c_{\boldsymbol{\nu}}\psi_{\boldsymbol{\nu}}.$ The high-dimensional polynomial basis $\{\psi_{\boldsymbol{\nu}}\}_{\boldsymbol{\nu}\in\mathbb{N}_{0}^{d}}$ is obtained as tensor product of such one-dimensional orthogonal polynomials $\{\phi_{k}\}_{k\in\mathbb{N}_{0}}$ over the interval [1,1][-1,1], i.e., $\psi_{\boldsymbol{\nu}}(\boldsymbol{x})=\prod_{k=1}^{d}\phi_{\nu_{k}}(x_{k})$, for $\boldsymbol{x}=(x_{1},\ldots,x_{d})\in\mathcal{U}$. Furthermore, we use the multi-index notation, i.e., instead of indexing a polynomial $\psi_{i}$ by $i\in\mathbb{N}$, we index it using a multi-index $\boldsymbol{\nu}=(\nu_{1},\nu_{2},...,\nu_{d})\in\mathbb{N}_{0}^{d}$. Typical examples of $\varrho$ and $\{\psi_{\boldsymbol{\nu}}\}_{\boldsymbol{\nu}\in\mathbb{N}_{0}^{d}}$ are the uniform measure and the multivariate Legendre polynomials, respectively, defined by
$\psi_{\boldsymbol{\nu}}(\boldsymbol{y})=\prod\limits_{k=1}^{d}\sqrt{2\nu_{k}+1}P_{\nu_{k}}(y_{k}),\quad\boldsymbol{y}\in[-1,1]^{d},\quad\forall\boldsymbol{\nu}\in\mathbb{N}_{0}^{d},$ (6)
where the $P_{\nu}$’s are the univariate Legendre polynomials defined, e.g., via Rodrigues’s formula: $P_{\nu}(y)=\frac{1}{2^{\nu}\nu!}\frac{d^{\nu}}{dy^{\nu}}(y^{2}-1)^{\nu}$, $y\in[-1,1]$, for all $\nu\in\mathbb{N}_{0}$. Another popular choice for $\varrho$ and $\{\psi_{\boldsymbol{\nu}}\}_{\boldsymbol{\nu}\in\mathbb{N}_{0}^{d}}$ is given by the Chebyshev measure and Chebyshev polynomials, respectively; see [1] for further details.
Our goal is to find an $s$-sparse polynomial approximation to $f$. In other words, we want $f\approx\sum_{\boldsymbol{\nu}\in S}c_{\boldsymbol{\nu}}\psi_{\boldsymbol{\nu}}$, where $S\subseteq\mathbb{N}_{0}^{d}$ and $\lvert S\rvert\leq s$.
@@
$\Lambda_{n}^{\mathrm{TD}}:=\bigg\{\boldsymbol{\nu}=\left(\nu_{k}\right)_{k=1}^{d}\in\mathbb{N}_{0}^{d}:\sum_{k=1}^{d}\nu_{k}\leq n\bigg\},$ (7)
@@
$\Lambda_{n}^{\mathrm{HC}}:=\bigg\{\boldsymbol{\nu}=\left(\nu_{k}\right)_{k=1}^{d}\in\mathbb{N}_{0}^{d}:\prod_{k=1}^{d}\left(\nu_{k}+1\right)\leq n+1\bigg\}.$ (8)
Note that the parameter $n\in\mathbb{N}_{0}$, called the order of the index set, and the dimension $d$ affect the cardinality of the index set. There is a closed-form expression for the cardinality of the total degree index set [1], given by $\lvert\Lambda_{n}^{\mathrm{TD}}\rvert={n+d\choose d}$, for all $d\in\mathbb{N}$, and $n\in\mathbb{N}_{0}$ The cardinality of the hyperbolic cross does not have an explicit formula, but it can be estimated as $\lvert\Lambda_{n-1}^{\mathrm{HC}}\rvert\leq\min\{2n^{3}4^{d},en^{2+\log_{2}(d)}\}$ (see [1, Appendix B] and references therein). In particular, $\lvert\Lambda_{n}^{\mathrm{HC}}\rvert\ll\lvert\Lambda_{n}^{\mathrm{TD}}\rvert$ for large $d$ or $n$. We illustrate these index sets in Figure 1, with $n=4$ and $d=3$.
Figure 1: Multi-index sets of order $n=4$ in $\mathbb{N}_{0}^{3}$ (left: total degree; right: hyperbolic cross).
@@
$\min_{\boldsymbol{z}\in\mathbb{R}^{s}}\lVert\boldsymbol{\Psi}\boldsymbol{z}-\boldsymbol{b}\rVert_{2}^{2}.$ (9)
An approximation $\boldsymbol{\hat{c}}$ to the vector $\boldsymbol{c}$ can be found using the Moore-Penrose pseudoinverse $\boldsymbol{\hat{c}}=(\boldsymbol{\Psi}^{*}\boldsymbol{\Psi})^{-1}\boldsymbol{\Psi}^{*}\boldsymbol{b}$. In order to retrieve a unique solution of a least-squares problem, we need an overdetermined system (a system with at least as many linearly independent equations as unknowns). In other words, for $\boldsymbol{\Psi}\in\mathbb{R}^{m\times s}$, we need $m\geq s$. This means that we need more sample points than coefficients $c_{\boldsymbol{\nu}_{1}},\dots,c_{\boldsymbol{\nu}_{s}}$ for this method to be effective. For further details on sparse polynomial approximation via least squares, we refer to [1, Chapter 5] and references therein.
On the other hand, suppose we have no a priori knowledge of $S$. In this case, a possible strategy is to choose a set $\Lambda\subseteq\mathbb{N}_{0}^{d}$ that contains $S$. A common choice of such $\Lambda$ in literature is represented by the hyperbolic cross, since it contains all so-called lower sets of a given cardinality; see Appendix A for further explanations). If we consider a larger set $\Lambda$ that contains $S$, we come to a point where the system is underdetermined. In other words, for $\boldsymbol{\Psi}\in\mathbb{C}^{m\times N}$, where $\Psi_{ij}=\frac{1}{\sqrt{m}}\psi_{\boldsymbol{\nu}_{j}}(\boldsymbol{x}_{i})$ for all $i=1,..,m$, j=1,,Nj=1,...,N, the cardinality $N=|\Lambda|$ of the set $\Lambda$ becomes greater than the number of sample points $m$. Therefore, if we want to recover the vector $\boldsymbol{c}\in\mathbb{C}^{N}$ of coefficients, this becomes a compressed sensing problem [24], where we attempt to find the sparsest solution to the system $\boldsymbol{\Psi}\boldsymbol{c}=\boldsymbol{b}$.
@@
$\min_{\boldsymbol{z}\in\mathbb{C}^{N}}\lambda\lVert\boldsymbol{z}\rVert_{1}+\lVert\boldsymbol{\Psi}\boldsymbol{z}-\boldsymbol{b}\rVert_{2}\text{, }$ (10)
@@
$\min_{\boldsymbol{z}\in\mathbb{C}^{N}}\lVert\boldsymbol{z}\rVert_{1}\quad\text{ subject to }\lVert\boldsymbol{\Psi}\boldsymbol{z}-\boldsymbol{b}\rVert_{2}\leq\eta\text{, }$ (11)
@@
To enhance the compressed sensing performance, a standard approach is to replace the $\ell^{1}$-norm by a weighted $\ell^{1}$-norm. This approach relies on such weights to represent a form of prior on the unknown vector to be recovered. Given a vector of weights $\boldsymbol{w}\in\mathbb{R}^{N}$ with $\boldsymbol{w}\geq\boldsymbol{0}$, we can replace the $\ell^{1}$-norm with $\|\boldsymbol{z}\|_{1,\boldsymbol{w}}:=\sum_{i=1}^{N}w_{i}z_{i}$ in equations (10) and (11). This yields, respectively, the weighted SR-LASSO problem
$\min\limits_{\boldsymbol{z}\in\mathbb{C}^{N}}\lambda\|\boldsymbol{z}\|_{1,\boldsymbol{w}}+\|\boldsymbol{\Psi}\boldsymbol{z}-\boldsymbol{b}\|_{2}$ (12)
and the weighted QCBP problem
$\min_{\boldsymbol{z}\in\mathbb{C}^{N}}\lVert\boldsymbol{z}\rVert_{1,\boldsymbol{w}}\quad\text{ subject to }\lVert\boldsymbol{\Psi}\boldsymbol{z}-\boldsymbol{b}\rVert_{2}\leq\eta.$ (13)
@@
High-dimensional functions may vary much faster in some coordinate directions than in others. This behavior is referred to as anisotropy. For such functions, choices of $S$ that treat all coordinate directions equally usually lead to poor approximations [1]. Typically, least squares methods are preferred for known anisotropy, while compressed sensing is the standard option for unknown anisotropy. We will be using both least-squares and compressed sensing in our numerical experiments, and compare their performance in different scenarios.
@@
$\sigma_{s}(\boldsymbol{c})_{p}:=\min\big\{\|\boldsymbol{c}-\boldsymbol{z}\|_{p}:\boldsymbol{z}\in\ell^{p}(\mathbb{N}_{0}^{d}),\;|\textnormal{supp}(\boldsymbol{z})|\leq s\big\}.$ (14)